OpenStudy (anonymous):

Mary will rent a car for the weekend. She can choose one of two plans. The first plan has an initial fee of $55.96 and costs an additional $0.14per mile driven. The second plan has an initial fee of $65.96 and costs an additional $0.12 per mile driven. How many miles would Mary need to drive for the two plans to cost the same?

OpenStudy (anonymous):

@kcbunny

OpenStudy (anonymous):

She would need to drive 500 miles for the two plants to cost the same

OpenStudy (anonymous):

I guess this is because...

OpenStudy (anonymous):

There is a $10 difference between the initial fees

OpenStudy (anonymous):

And a 2 cents difference between additional miles driven

OpenStudy (anonymous):

Thanks! do you know how to do this one? There are 240 pounds on the left plate of a scale and 90 pounds on the right plate. Suppose that you must balance the scale by removing 8-pound bags from the left plate and exactly the same number of 3-pound bags from the right plate. How many bags must you remove from each side in order to balance the scale?

OpenStudy (anonymous):

Are you familiar with how to solve two variable equations?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Actually, I'm not sure you need it. Hang on. Sorry

OpenStudy (anonymous):

@justin12lo, if you are still here and want to help with this, that's fine. I know there is a really easy way to figure this out but right now I'm stuck on just subtracting 80 from one side and adding 30 to the other until they are even

OpenStudy (anonymous):

8 and 3* sorry

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

I'm stuck, the only way I see it is by removing all the bags. It's confusing me at the moment :(

OpenStudy (anonymous):

Glad to see we are in the same boat. We'll get this @ashleynicolee :)

OpenStudy (anonymous):

There can't be nothing

OpenStudy (anonymous):

haha thanks guys

OpenStudy (anonymous):

I think it's going to be a matter of finding a common factor. Really you have 0 on one plate and 150 on the other

OpenStudy (anonymous):

I'll come back in a few minutes if I have a solution

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Otherwise I will keep trying

OpenStudy (anonymous):

could it be possible you just have to remove all the bags?

OpenStudy (anonymous):

That's what I was thinking at first...

OpenStudy (anonymous):

I don't know if there is a trick or something or it really is remove every bag.

OpenStudy (anonymous):

seems odd that you would have to do that. I suppose it is possible, but what if there aren't any bags on the 90 pound side and thats just the weight of the scale?

OpenStudy (anonymous):

i tried that. and its wrong :/

OpenStudy (anonymous):

I'm calling my dad. He has a Masters in math so I'll just ask him

OpenStudy (anonymous):

well thats extremely helpful!

OpenStudy (anonymous):

okay!! I knew there was a super simple way. Okay, so you set it up in an equation. Because you want both sides equal, they are set equal to each other: Left side = right side At the beginning of the problem, we have 240 on one side and 90 on the other: 240 = 90 (I know it really doesn't, but bear with me as I set it up) Next step is that we are removing 8 pounds from the left side and 3 pounds from the right: 240 - 8 = 90 - 3 Then we know that you have to remove the same number of bags (let x stand for the number of bags) from both sides and that number is multiplied by the number of pounds: 240 - 8x = 90 - 3x Solve for x

OpenStudy (anonymous):

so the answer is 30

OpenStudy (anonymous):

Yes. You have to remove 30 bags which is actually all of them.

OpenStudy (anonymous):

I totally misread the question the first time and thought we were adding 3 pounds on the right side while taking off 8 on the left.